OpenStudy (anonymous):

Solve quadratic equation: x+2/x+1 = x+2/1-x - 4/x-1 Domain is - x1 =/= 1 ; x2 =/= -1 Help me solve it? In my country we have to multiply the numerators with the greatest common factors (that are from the denominators) and reach to a solution with x1 and x2 with the quadratic equation formula -b +- sqrt b^2 - 4ac/ 2a. Please help!!

OpenStudy (anonymous):

\[\frac{ x+2 }{ x+1 } = \frac{ x+2 }{ 1-x} - \frac{ 4 }{ x-1 }\]

OpenStudy (anonymous):

\[\(\bf\huge\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\]

OpenStudy (anonymous):

thank you. do you know how to solve this? :/

OpenStudy (anonymous):

(1-x) 0r (x-1) is the same @galaxy-fanfan

OpenStudy (anonymous):

\[\frac{ x+2 }{ x-1 }=\frac{ x+2-4 }{ x-1 }\]

OpenStudy (anonymous):

\[\frac{ x+2 }{ x-1 }=\frac{ x-2 }{ x-1 }\]

OpenStudy (anonymous):

do cross method @galaxy-fanfan

OpenStudy (triciaal):

one approach first multiply x +2/1-x by -1/-1 this will change the denominator to (x -1) and the numerator to -x -2 rewrite the original with this

OpenStudy (anonymous):

thnx @triciaal

OpenStudy (triciaal):